OpenStudy (anonymous):

Simplify the complex fraction. (2-2/x)/(3-1/x)

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[\frac{2-\frac{2}{x}}{3-\frac{1}{x}}\] like that?

OpenStudy (anonymous):

Yes exactly like that!

OpenStudy (isaiah.feynman):

Evaluate top and bottom.

OpenStudy (anonymous):

I am so confused i don't even know where to start

OpenStudy (isaiah.feynman):

Multiply top and bottom by x.